# DRAN: A Distribution and Relation Adaptive Network for Spatio-temporal Forecasting

### v2

@@ III-A Problem Definition
5:        Obtain $\bm{\textit{X}}_{\rm Norm}$ via Eq (1)
6:        $\bm{\textit{X}}_{\rm tem}\xleftarrow{}\rm{DestationaryAtt}(\bm{\textit{X}}_{\rm Norm},\mu_{\bm{X}},\sigma_{\bm{X}})$ via Eq (2-4)
@@
8:        Rescale according to Eq (5) and obtain $\bm{\textit{X}}_{\rm spa}$
9:        $\rm{\textit{X}}_{\rm D}\xleftarrow{}\rm{DSFL}(\bm{\textit{X}}_{\rm spa},\bm{\textit{E}}_{\rm a})$ via Eq (6-11)
10:        $\rm{\textit{X}}_{\rm S},\rm{\textit{X}}_{\rm rec}\xleftarrow{}\rm{StochasticLearner}(\bm{\textit{X}}_{\rm D})$ via Eq (12-15)
11:        Compute $\mathcal{L}$ via Eq (17) and optimize $\theta$
@@ III-C Dynamic-Static Fusion Learner
where $\rm F_{\text{Cat}}$ conducts concatenation operation on the last dimension of dynamic features $\bm{X}_{\rm Dy}$ and static features $\bm{X}_{\rm St}$, generating the concatenated features $\bm{X}_{\rm Cat}\in\mathbbm{R}^{L\times N\times 2C}$. We generate the gate signal $\bm{z}$ by applying a FC layer to map $\bm{X}_{\rm Cat}$ from $\mathbbm{R}^{L\times N\times 2C}$ to $\mathbbm{R}^{L\times N\times C}$, followed by a Sigmoid activation function; here, FC denotes the fully connected layer. Finally, we employ Equation (12) to fuse the dynamic and static features, producing the final deterministic features $\bm{X}_{\rm D}$, which represent the final deterministic features learned from the input time series.
@@ V-C Dynamic and Static Relations Learning
In Fig. 7, we explore the adaptive process of our model by depicting the dynamic and static adjacency matrices within the DSFL module. Specifically, we showcase these matrices for the first input time step as an example. The dynamic adjacency matrix $\bm{A}_{\rm Dy}$ is derived from the similarity of time series between nodes according to Equation 6, while the static adjacency matrix is obtained by learning the static relations $\bm{A}_{\rm St}$ according to Equation 8. Fig. 7 illustrates the spatial relations between nodes of Weather dataset, with darker colors indicating stronger relationships. The dynamic adjacency matrix highlights the strength of relationships between nodes with similar signal patterns, whereas the static adjacency matrix focuses on the signals of individual nodes and some distributed nodes within the network. The contrast in focus enables the model to learn complementary features from both dynamic and static perspectives.

### v3

@@ III-A Problem Definition
5:        Obtain $\bm{\textit{X}}_{\rm Norm}$ via Eq. (1)
6:        $\bm{\textit{X}}_{\rm tem}\xleftarrow{}\rm{DestationaryAtt}(\bm{\textit{X}}_{\rm Norm},\mu_{\bm{X}},\sigma_{\bm{X}})$ via Eq. (2-4)
@@
8:        Rescale according to Eq. (5) and obtain $\bm{\textit{X}}_{\rm spa}$
9:        $\rm{\textit{X}}_{\rm D}\xleftarrow{}\rm{DSFL}(\bm{\textit{X}}_{\rm spa},\bm{\textit{E}}_{\rm a})$ via Eq. (6-11)
10:        $\rm{\textit{X}}_{\rm S},\rm{\textit{X}}_{\rm rec}\xleftarrow{}\rm{StochasticLearner}(\bm{\textit{X}}_{\rm D})$ via Eq. (12-15)
11:        Compute $\mathcal{L}$ via Eq. (17) and optimize $\theta$
@@ III-C Dynamic-Static Fusion Learner
where $\rm F_{\text{Cat}}$ conducts concatenation operation on the last dimension of dynamic features $\bm{X}_{\rm Dy}$ and static features $\bm{X}_{\rm St}$, generating the concatenated features $\bm{X}_{\rm Cat}\in\mathbbm{R}^{L\times N\times 2C}$. We generate the gate signal $\bm{z}$ by applying a FC layer to map $\bm{X}_{\rm Cat}$ from $\mathbbm{R}^{L\times N\times 2C}$ to $\mathbbm{R}^{L\times N\times C}$, followed by a Sigmoid activation function; here, FC denotes the fully connected layer. Finally, we employ Eq. (12) to fuse the dynamic and static features, producing the final deterministic features $\bm{X}_{\rm D}$, which represent the final deterministic features learned from the input time series.
@@ V-C Dynamic and Static Relations Learning
In Fig. 7, we explore the adaptive process of our model by depicting the dynamic and static adjacency matrices within the DSFL module. Specifically, we showcase these matrices for the first input time step as an example. The dynamic adjacency matrix $\bm{A}_{\rm Dy}$ is derived from the similarity of time series between nodes according to Eq. 6, while the static adjacency matrix is obtained by learning the static relations $\bm{A}_{\rm St}$ according to Eq. 8. Fig. 7 illustrates the spatial relations between nodes of Weather dataset, with darker colors indicating stronger relationships. The dynamic adjacency matrix highlights the strength of relationships between nodes with similar signal patterns, whereas the static adjacency matrix focuses on the signals of individual nodes and some distributed nodes within the network. The contrast in focus enables the model to learn complementary features from both dynamic and static perspectives.